OpenStudy (anonymous):

How do you write the quadratic function f(x) = 3x2 - 6x + 9 in vertex form

OpenStudy (anonymous):

3(x^2 - 2x + 3) = 3 [(x-1)^2 -1 + 3] = 3(x-1)^2 + 6

OpenStudy (anonymous):

we just did this in school... the equation for vertex form is: \[y=a(x-h)^{2}+k\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome (: so if you were doing the equation you posted, you would change the problem to y=3x2-6x+9 because f(x) is another way to say y. do you get it now? or do you need help with the problem?

OpenStudy (anonymous):

i think you'll find i've done it correctly

OpenStudy (anonymous):

i understand, but i thought it nice to explain.

OpenStudy (anonymous):

yup - thats ok